Accelerator Update for January 11, 2008 - January 14, 2008
Friday January 11
The day shift began with Operations monitoring the Tevatron (TeV) store 5838 (116.17E30) and the Recycler stash (138.2E10), with the Antiproton Source stacking (11.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 8:17 AM, KRF4 tripped off. It reset.
At 8:18 AM, LRF2 tripped off.
An Instrumentation expert reported at 9:59 AM that he was updating the code for the MI flying wires.
At 10:38 AM, KRF7 tripped off. It reset.
At 12:42 PM, CDF reported that they were turning off their solenoid due to a cryo system problem.
Operations reported around 1 PM that they had turned off the Booster beam so technicians could repair BRF15. Booster resumed sending beam around 1:44 PM.
At 2:10 PM, LRF2 tripped off. It reset.
The Neutron Therapy Facility (NTF) reported at 2:26 PM that they had completed treating their last patient for the day.
Operations reported at 3:54 PM that the Recycler had stashed antiprotons five times during the day shift.
At 5:03 PM, KRF5 tripped off. It reset.
CDF reported at 7:52 PM that they were turning on the solenoid. They resumed taking data at 9:03 PM.
At 11:12 PM, MIRF2 tripped off and would not reset. Operators contacted an expert and then bypassed the station. Operations resumed sending beam to NuMI at 11:34 PM.
Operations reported at midnight that the Recycler had stashed five times during the evening shift.
Saturday January 12
The midnight shift began with Operations monitoring store 5838 and the stash (399E10), with Pbar stacking (24.4E10), and with NuMI and BNB taking beam.
Operations terminated store 5838 at 12:11 AM.
Operations put the TeV into shot setup at 12:13 AM.
At 1:08 AM, LRF2 tripped off for the first of three times in about two hours. It reset.
At 1:28 AM, Operations reported that due to problems with a Pbar quadrupole power supply (I:Q701) the P1 transfer line was tripping off and taking down NuMI with it. Operations called in an expert.
Operations reported at 4:08 AM, that operators changed the I:Q701 ramp card.
Operations restarted the shot setup at 4:50 AM; this was due to the long time that the TeV was sitting at 150 GeV.
At 5:05 AM, LRF2 tripped off. It reset.
At 5:17 AM, the Booster West Anode Power Supply (WAPS) tripped off and would not reset. Operations investigated and found that BRF18 was the problem. Operators bypassed the station at 5:47 AM and Booster resumed sending beam.
At 6:02 AM, Operations reported that they were having problems with reverse injections into the TeV and contacted an expert.
At 7:03 AM, Operations reported that I:V714 was acting as I:Q701 had acted. Operators changed its ramp card, but this didn’t fix the problem. Operators contacted a Controls expert.
An expert reported he had repaired the load switch for I:V714 and had it working properly by 9:21 AM.
Operations put the TeV into shot setup at 9:57 AM.
Operations established store 5841 at 12:35 PM with an initial luminosity of 221.61E30.
At 1:45 PM, a NuMI power supply (E:V108) tripped off on a ground fault for the first of two times in a minute. It reset.
At 2:16 PM, the TeV RF watchdog alarmed on a pond water pressure indications. Operations contacted the Duty Mechanic and he reported that the strainer was beginning to clog. The RF expert said that this would be okay until next week when divers would clean the strainer.
At 3:36 PM, LRF2 tripped off. It reset.
At 5:24 PM, Operations reported having MI beam position problems in the 522 area. Operators worked on the problem and then called in an expert.
Operations reported at 8:02 PM that the Recycler had stashed five times during the last 12 hours.
At 8:09 PM, LRF2 tripped off. It reset.
A MI expert completed tuning MI at 9:21 PM.
The MI expert who was tuning requested Operations at 11:05 PM take the cavity short off of MIRF2. This made MI beam much more stable.
At 11:27 PM, a Booster kicker (MKS08) tripped off and would not reset. Operators adjusted the other kickers to take over the load and then called in experts to make repairs.
Sunday January 13
The midnight shift began with Operations monitoring store 5841 and the stash, with Pbar stacking, and with NuMI and BNB taking beam.
At 12:24 AM, MIRF15 tripped off. It reset.
Booster experts reported at 12:50 AM that they had repaired a PFN cable for MKS08.
At 1:20 AM, LRF2 tripped off and would not reset. Operators called in Linac experts.
At 6:20 AM, a TeV RF station (TRF1) tripped off on an upstream heater ground fault. Operators contacted an RF expert.
A Linac expert reported at 7:01 AM that LRF2 was repaired.
Operations reestablished beam to BNB, NuMI, and Pbar by 7:02 AM.
At 7:22 AM, LRF2 tripped off twice in 21 minutes. It reset.
Operations reported at 7:52 AM that the Recycler had stashed three times during the last 12 hours.
At 8:32 AM, LRF2 tripped off. It reset.
Operations terminated store 5841 at 9:08 AM.
Operators turned TRF1 back on at 9:09 AM.
Operations put the TeV into shot setup at 9:43 AM.
Operations established store 5842 at 11:53 AM with an initial luminosity of 209.02E30.
Operations reported at 8 PM that the Recycler had stashed six times during the last 12 hours.
At 9:31 PM, LRF2 tripped off for the first of two times in about two hours. It reset.
Monday January 14
The midnight shift began with Operations monitoring store 5842 and the stash, with Pbar stacking, and with NuMI and BNB taking beam.
At 1:01 AM, the Pbar beam permit began tripping off. The stack remained intact, but every Pbar trip was causing NuMI to trip off. Everything reset.
At 2:19 AM, KRF3 tripped off. It reset.
Operations reported at 8:02 AM that the Recycler had stashed eight times during the last 12 hours.
The Plan for Monday
The plan for today is to stack, stash, and store. The Run Coordinator will allow Booster experts to conduct a one-hour damper study.
Comments and Suggestions
|last modified 01/14/2008 email Fermilab|
|Security, Privacy, Legal|